• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 14:25:59 on February25,2002

    Entry number 4171

    keyword=DAQ problems, fixed, we think

    There were three observed problems with DAQ: 1) Since Friday we had noticed
    that running test code on the ADC in slot 25 of roc2 led to SFI arbitration errors;
    2) Large fluctuations in roc2 busy signal - so we had removed this crate from the
    OR'd busy, which was dangerous. It was noted Friday. 3) Today the experiment
    complained about deadtime fluctuating from 20 to 80 %.
    ----
    We took access and replaced the ADC 1881 in slot 25. This solved the first 2
    problems and probably the 3rd, though its hard to say about the 3rd. I ran DAQ
    at 1.8 kHz, deadtime stable. The roc2 busy is plugged into the OR'd busy again.